# Generalized Linear Bandits: Almost Optimal Regret with One-Pass Update

### v1

@@ Appendix B Proof of Theorem 2
[…]
@@
[…]
@@
[…]
where the first inequality is due the self-concordant property of $\mu$. The second inequality is by Assumption 2 such that $\mu^{\prime}(z)\leq C_{\mu}$ for $z\in[-S,S]$ and the Hölder’s inequality. The last inequality is due to Theorem 1 and the fact $H_{t+1}\succcurlyeq H_{t}$.
@@
$\mathtt{term~{}(a1)}\leq{}$ $\frac{2\beta_{T}(\delta)}{\sqrt{g(\tau)}}\sum_{t\in\mathcal{T}_{1}}\mu^{\prime}(X_{t}^{\top}\theta_{t+1})\lVert X_{t}\rVert_{\widetilde{H}_{t}^{-1}}+\frac{2RC_{\mu}\beta^{2}_{T+1}(\delta)}{g(\tau)}\sum_{t=1}^{T}\lVert X_{t}\rVert^{2}_{\widetilde{H}_{t}^{-1}}$
[…]
@@
$\sum_{t\in\mathcal{T}_{1}}\left\lVert\sqrt{{\mu^{\prime}(X_{t}^{\top}\theta_{t+1})}}X_{t}\right\rVert_{\widetilde{H}_{t}^{-1}}^{2}\leq{}$ $\sum_{t\in[T]}\left\lVert\sqrt{{\mu^{\prime}(X_{t}^{\top}\theta_{t+1})}}X_{t}\right\rVert_{\widetilde{H}_{t}^{-1}}^{2}\leq 2d(1+C_{\mu})\log\left(1+\frac{TC_{\mu}}{g(\tau)d\lambda}\right)$ (31)
@@
$\begin{cases}\gamma_{T}^{(1)}(\delta)=\frac{2\beta_{T}(\delta)\sqrt{2d(1+C_{\mu})}}{\sqrt{g(\tau)}}\sqrt{\log\left(1+\frac{TC_{\mu}}{g(\tau)d\lambda}\right)}=\mathcal{O}(\beta_{T}(\delta)\sqrt{d\log T}),\\ \gamma_{T}^{(2)}(\delta)=\frac{8\kappa dRC_{\mu}\beta^{2}_{T+1}(\delta)}{g(\tau)}\log\bigg{(}1+\frac{T}{g(\tau)\kappa d\lambda}\bigg{)}=\mathcal{O}\left(\beta_{T+1}(\delta)^{2}\kappa dR\log T\right).\end{cases}$ (33)
@@
[…]
@@
As shown in Proposition 1 in Appendix D, the time complexity for updating the online estimator $\theta_{t}$ is $\mathcal{O}(d^{3})$ (line 5 in Algorithm 1). Additionally, the inverse Hessian matrix $H_{t}^{-1}$ can be updated in $\mathcal{O}(d^{2})$ time per round as shown in Appendix D. The remaining computational cost arises from the arm selection (6), which solves the optimization problem
@@ Appendix C Technical Lemmas
Lemma 9 (Theorem 3.9 of Lattimore and Szepesvári [2020]).
Let $\{M_{t}\}_{t=0}^{\infty}$ be a supermargingle with $M_{t}\geq 0$ almost surely for all $t\geq 0$. Then, for any $\varepsilon>0$,
@@
Let $P=\mathcal{N}(\mathbf{0},\eta H^{-1})$ be a Gaussian distribution with mean $\mathbf{0}\in\mathbb{R}^{d}$ and covariance $\eta H^{-1}$, where $H\succcurlyeq\lambda I_{d}$ is a positive definite matrix and $\eta>0$ Then, if $\lambda\geq 32d\eta c/7$ we have
@@
[…]
@@
[…]
@@
Appendix D Computational Cost Discussion
This section discusses the time and space complexity of solving the optimization problem (3).
Proposition 1.
Space Complexity. Regarding space complexity, it suffices to store the current model $\theta_{t}$, the gradient $\nabla\ell_{t}(\theta_{t})$, the inverse Hessian matrix $H_{t}^{-1}$, and $\widetilde{H}_{t}^{-1}$ throughout the optimization process, resulting in a total space complexity of $\mathcal{O}(d^{2})$.
Appendix E Detailed Discussion about Lee and Oh [2025]
@@
[…]
where ¯$\bar{\ell}_{s}(z)=r_{s}z+\log(1+\exp(z))$. In Lee and Oh [2025], the intermediate term is chosen as $\widetilde{z}_{s}=\sigma^{+}\left(\mathbb{E}_{\theta\sim{P_{s}}}[\sigma(X_{s}^{\top}\theta)]\right)$, where $\sigma(z)=1/(1+e^{-z})$ and $\sigma^{+}(p)=\log(\frac{p}{1-p})$.
@@
Appendix F Additional Experimental Details
@@ F.1 Experimental Setup
All the experiments were conducted on Intel Xeon Gold 6242R processors (40 cores, 4.1GHz base frequency). The algorithms were implemented in Python, utilizing the scipy library for numerical computations, such as solving non-linear optimization problems and calculating vector norms, and employing np.linalg.pinv to compute the pseudo-inverse of matrices.
@@
F.2 More results on Synthetic Data
@@
To further investigate the impact of parameter $S$ on algorithm performance, we conduct additional experiments on logistic bandit tasks with larger $S$ values (Figures 3(a) and 3(b)) and analyze the computational time scaling for Poisson bandits (Figure 3(c)).
The regret curves in Figures 3(a) and 3(b) consistently demonstrate that our algorithm maintains its competitive performance regardless of $S$ variations, aligning with the trends observed in our main results. Notably, the regret does not exhibit significant growth as $S$ increases, suggesting the robustness of our approach to parameter scaling. We also note that the performance of RS-GLinCB is very sensitive to the parameter of $S$. This underperformance can be attributed to the fact that the warm-up period of RS-GLinCB is heavily dependent on the constant $S$ and $\kappa$ [Sawarni et al., 2024, Lemma 4.1].
@@
where $\beta_{t}(\delta)^{2}=\log\frac{1}{\delta}+\inf_{c_{t}\in(0,1]}\left\{d\log\frac{1}{c_{t}}+2SL_{t}c_{t}\right\}\leq\log\frac{1}{\delta}+d\log\left(e\vee\frac{2eSL_{t}}{d}\right)$. For Poisson bandits specifically, $L_{t}=e^{S}t+\log(d/\delta)$. This results in increasing cost in the optimization steps during maximum likelihood estimation, as the algorithm must navigate a rapidly expanding confidence region.
@@
F.3 Experiment on Forest Cover Type Data
@@
For this task, we set the horizon to $T=1000$ and the confidence parameter to $\delta=0.01$. After analyzing the data, we set $S=6$ and $\kappa=200$. We evaluated our algorithm against the same baselines used in the logistic bandit simulation experiment, running each method over 10 independent trials and averaging the results to report the regret and the running time. Compared to synthetic data, real-world datasets exhibit higher noise and complexity, demanding careful exploration-exploitation trade-offs. Thus, we shrank the estimated confidence set of all the algorithms in a comparable way to achieve a better balance between exploration and exploitation in this real-world dataset. Traditional GLB algorithms are particularly sensitive to noise, often leading to excessive exploration and higher regret.

### v2

@@ Appendix B Proof of Theorem 2
$\leq{}$ $\sum_{t=1}^{T}\mu(X_{t}^{\top}\widetilde{\theta}_{t})-\sum_{t=1}^{T}\mu(X_{t}^{\top}\theta_{*})$
$={}$ $\underbrace{\sum_{t=1}^{T}\mu^{\prime}(X_{t}^{\top}\theta_{*})X_{t}^{\top}(\widetilde{\theta}_{t}-\theta_{*})}_{\mathtt{term~(a)}}+\underbrace{\frac{1}{2}\sum_{t=1}^{T}\widetilde{\alpha}(\theta_{*},\widetilde{\theta_{t}},X_{t})\big(X_{t}^{\top}(\widetilde{\theta}_{t}-\theta_{*})\big)^{2}}_{\mathtt{term~(b)}},$ (29)
@@
$\leq{}$ $\sum_{t=1}^{T}\mu^{\prime}(X_{t}^{\top}\theta_{*})\cdot\lVert X_{t}\rVert_{H_{t}^{-1}}\lVert\widetilde{\theta}_{t}-\theta_{*}\rVert_{H_{t}}$
$\leq{}$ $2\sum_{t=1}^{T}\mu^{\prime}(X_{t}^{\top}\theta_{*})\cdot\beta_{t}(\delta)\lVert X_{t}\rVert_{H_{t}^{-1}}$
$\leq{}$ $2\beta_{T}(\delta)\sum_{t=1}^{T}\mu^{\prime}(X_{t}^{\top}\theta_{*})\cdot\lVert X_{t}\rVert_{H_{t}^{-1}}$
$\leq{}$ $\underbrace{2\beta_{T}(\delta)\sum_{t\in\mathcal{T}_{1}}\mu^{\prime}(X_{t}^{\top}\theta_{*})\cdot\lVert X_{t}\rVert_{H_{t}^{-1}}}_{\mathtt{term~(a1)}}+\underbrace{2\beta_{T}(\delta)\sum_{t\in\mathcal{T}_{2}}\mu^{\prime}(X_{t}^{\top}\theta_{*})\cdot\lVert X_{t}\rVert_{H_{t}^{-1}}}_{\mathtt{term~(a2)}},$
@@
$\leq{}$ $\mu^{\prime}(X_{t}^{\top}\theta_{t+1})+R\int_{v=0}^{1}\mu^{\prime}\big(X_{t}^{\top}\theta_{t+1}+vX_{t}^{\top}(\theta_{*}-\theta_{t+1})\big)\mathrm{d}v\cdot|X_{t}^{\top}(\theta_{*}-\theta_{t+1})|$
$\leq{}$ $\mu^{\prime}(X_{t}^{\top}\theta_{t+1})+RL_{\mu}\cdot\lVert X_{t}\rVert_{H_{t+1}^{-1}}\cdot\lVert\theta_{*}-\theta_{t+1}\rVert_{H_{t+1}}$
$\leq{}$ $\mu^{\prime}(X_{t}^{\top}\theta_{t+1})+RL_{\mu}\beta_{t+1}(\delta)\cdot\lVert X_{t}\rVert_{H_{t}^{-1}}$ (30)
where the first inequality is due the self-concordant property of $\mu$. The second inequality is by Assumption 2 such that $\mu^{\prime}(z)\leq L_{\mu}$ for $z\in[-S,S]$ and the Hölder’s inequality. The last inequality is due to Theorem 1 and the fact $H_{t+1}\succcurlyeq H_{t}$.
@@
$\mathtt{term~(a1)}\leq{}$ $\frac{2\beta_{T}(\delta)}{\sqrt{g(\tau)}}\sum_{t\in\mathcal{T}_{1}}\mu^{\prime}(X_{t}^{\top}\theta_{t+1})\lVert X_{t}\rVert_{\widetilde{H}_{t}^{-1}}+\frac{2RL_{\mu}\beta^{2}_{T+1}(\delta)}{g(\tau)}\sum_{t=1}^{T}\lVert X_{t}\rVert^{2}_{\widetilde{H}_{t}^{-1}}$
$\leq{}$ $\frac{2\beta_{T}(\delta)}{\sqrt{g(\tau)}}\sum_{t\in\mathcal{T}_{1}}\sqrt{{\mu^{\prime}(X_{t}^{\top}\theta_{*})}}\cdot\left\lVert\sqrt{\mu^{\prime}(X_{t}^{\top}\theta_{t+1})}X_{t}\right\rVert_{\widetilde{H}_{t}^{-1}}+\frac{2RL_{\mu}\beta^{2}_{T+1}(\delta)}{g(\tau)}\sum_{t\in\mathcal{T}_{1}}\lVert X_{t}\rVert^{2}_{\widetilde{H}_{t}^{-1}}$
$\leq{}$ $\frac{2\beta_{T}(\delta)}{\sqrt{g(\tau)}}\sqrt{\sum_{t\in\mathcal{T}_{1}}\mu^{\prime}(X_{t}^{\top}\theta_{*})}\cdot\sqrt{\sum_{t\in\mathcal{T}_{1}}\left\lVert\sqrt{{\mu^{\prime}(X_{t}^{\top}\theta_{t+1})}}X_{t}\right\rVert_{\widetilde{H}_{t}^{-1}}^{2}}+\frac{2RL_{\mu}\beta^{2}_{T+1}(\delta)}{g(\tau)}\sum_{t\in\mathcal{T}_{1}}\lVert X_{t}\rVert^{2}_{V_{t}^{-1}},$ (31)
@@
$\sum_{t\in\mathcal{T}_{1}}\left\lVert\sqrt{{\mu^{\prime}(X_{t}^{\top}\theta_{t+1})}}X_{t}\right\rVert_{\widetilde{H}_{t}^{-1}}^{2}\leq{}$ $\sum_{t\in[T]}\left\lVert\sqrt{{\mu^{\prime}(X_{t}^{\top}\theta_{t+1})}}X_{t}\right\rVert_{\widetilde{H}_{t}^{-1}}^{2}\leq 2d(1+L_{\mu})\log\left(1+\frac{TL_{\mu}}{g(\tau)d\lambda}\right)$ (32)
@@
$\begin{cases}\gamma_{T}^{(1)}(\delta)=\frac{2\beta_{T}(\delta)\sqrt{2d(1+L_{\mu})}}{\sqrt{g(\tau)}}\sqrt{\log\left(1+\frac{TL_{\mu}}{g(\tau)d\lambda}\right)}=\mathcal{O}(\beta_{T}(\delta)\sqrt{d\log T}),\\ \gamma_{T}^{(2)}(\delta)=\frac{8\kappa dRL_{\mu}\beta^{2}_{T+1}(\delta)}{g(\tau)}\log\bigg(1+\frac{T}{g(\tau)\kappa d\lambda}\bigg)=\mathcal{O}\left(\beta_{T+1}(\delta)^{2}\kappa dR\log T\right).\end{cases}$ (34)
@@
$\leq{}$ $\frac{R}{2}\sum_{t=1}^{T}\int_{0}^{1}\mu^{\prime}\big(X_{s}^{\top}\theta_{*}+v\,X_{s}^{\top}(\widetilde{\theta}_{t}-\theta_{*})\big)\,\mathrm{d}v\big(X_{t}^{\top}(\widetilde{\theta}_{t}-\theta_{*})\big)^{2}$
$\leq{}$ $\frac{RL_{\mu}}{2}\sum_{t=1}^{T}\lVert\widetilde{\theta}_{t}-\theta_{*}\rVert_{H_{t}}^{2}\cdot\lVert X_{t}\rVert_{H_{t}^{-1}}^{2}$
$\leq{}$ $\frac{2RL_{\mu}\beta^{2}_{T}(\delta)}{g(\tau)}\sum_{t=1}^{T}\lVert X_{t}\rVert_{V_{t}^{-1}}^{2}\leq\gamma_{T}^{(2)}(\delta),$~ (36)
@@
As shown in Proposition 1 in Appendix A.3, the time complexity for updating the online estimator $\theta_{t}$ is $\mathcal{O}(d^{3})$ (line 5 in Algorithm 1). Additionally, the inverse Hessian matrix $H_{t}^{-1}$ can be updated in $\mathcal{O}(d^{2})$ time per round as shown in Appendix A.3. The remaining computational cost arises from the arm selection (6), which solves the optimization problem
@@ Appendix C Technical Lemmas
Lemma 9 (Ville [1939]).
Let $\{M_{t}\}_{t=0}^{\infty}$ be a supermartingale with $M_{t}\geq 0$ almost surely for all $t\geq 0$. Then, for any $\varepsilon>0$,
@@
Let $P=\mathcal{N}(\mathbf{0},\eta H^{-1})$ be a Gaussian distribution with mean $\mathbf{0}\in\mathbb{R}^{d}$ and covariance $\eta H^{-1}$, where $H\succcurlyeq\lambda I_{d}$ is a positive definite matrix and $\eta>0$ . Then, if $\lambda\geq 32d\eta c/7$ we have
@@
$={}$ $\bigg(\mathbb{E}_{Z\sim\chi^{2}}\Big[\exp\Big(\frac{c\eta}{\lambda}Z\Big)\Big]\bigg)^{d}\leq\mathbb{E}_{Z\sim\chi^{2}}\Big[\exp\Big(\frac{c\eta d}{\lambda}Z\Big)\Big]\leq\frac{4}{3}$
@@
$={}$ $\frac{1}{\alpha}\mathbb{E}_{\theta\sim Q}\left[\ln\Big(\frac{P_{*}(\theta)}{P(\theta)}\Big)\right]-\frac{1}{\alpha}\mathbb{E}_{\theta\sim P_{*}}\left[\ln\Big(\frac{P_{*}(\theta)}{P(\theta)}\Big)\right]$
$={}$ $\frac{1}{\alpha}\mathbb{E}_{\theta\sim Q}\left[\ln\Big(\frac{e^{-\alpha\ell(\theta)}}{Z}\Big)\right]-\frac{1}{\alpha}\mathbb{E}_{\theta\sim P_{*}}\left[\ln\Big(\frac{e^{-\alpha\ell(\theta)}}{Z}\Big)\right]$
$={}$ $\mathbb{E}_{\theta\sim P_{*}}[\ell(\theta)]-\mathbb{E}_{\theta\sim Q}[\ell(\theta)].$
@@
Appendix D More Discussions on Lee and Oh [2025]
@@
$\lVert\theta_{t+1}-\theta_{*}\rVert^{2}_{H_{t+1}}\lesssim\underbrace{\sum_{s=1}^{t}\ell_{s}(\theta_{*})-\sum_{s=1}^{t}\bar{\ell}_{s}(\widetilde{z}_{s})}_{\texttt{term~(a)}}+\underbrace{\sum_{s=1}^{t}\bar{\ell}_{s}(\widetilde{z}_{s})-\sum_{s=1}^{t}\ell_{s}(\theta_{s+1})}_{\texttt{term~(b)}},$
where $\bar{\ell}_{s}(z)=r_{s}z+\log(1+\exp(z))$. In Lee and Oh [2025], the intermediate term is chosen as $\widetilde{z}_{s}=\sigma^{+}\left(\mathbb{E}_{\theta\sim{P_{s}}}[\sigma(X_{s}^{\top}\theta)]\right)$, where $\sigma(z)=1/(1+e^{-z})$ and $\sigma^{+}(p)=\log(\frac{p}{1-p})$.
@@
Appendix E Additional Experimental Details
@@ E.1 Experimental Setup
All the experiments were conducted on Intel Xeon Gold 6242R processors (40 cores, 4.1GHz base frequency). The algorithms were implemented in Python, utilizing the scipy library for numerical computations, such as solving non-linear optimization problems and calculating vector norms, and employing np.linalg.pinv to compute the pseudo-inverse of matrices. The running time was measured using the time library. The shaded regions in the regret plots represent 99% confidence intervals, computed from 10 independent runs with different random seeds.
@@
E.2 More results on Synthetic Data
(a) $d=2,S=3,T=300$.
(b) $d=2,S=5,T=300$.
(c) $d=2,S=7,T=300$.
Figure 3: Confidence Region of Parameter Estimation.
@@ E.2 More results on Synthetic Data
To visualize the accuracy of parameter estimation of the algorithms, we plot the confidence region of the parameter estimation for each algorithm in Figure 3. For illustration purposes, we only plot the confidence regions of our algorithm GLB-OMD, the theoretically optimal OFUGLB, and the classical GLM-UCB. We observe that both GLB-OMD and OFUGLB achieve a substantially smaller confidence region than GLM-UCB, indicating that our algorithm achieves an accurate parameter estimation comparable to the statistically optimal.
To further investigate the impact of parameter $S$ on algorithm performance, we conduct additional experiments on logistic bandit tasks with larger $S$ values (Figures 4(a) and 4(b)) and analyze the computational time scaling for Poisson bandits (Figure 4(c)).
The regret curves in Figures 4(a) and 4(b) consistently demonstrate that our algorithm maintains its competitive performance regardless of $S$ variations, aligning with the trends observed in our main results. Notably, the regret does not exhibit significant growth as $S$ increases, suggesting the robustness of our approach to parameter scaling. We also note that the performance of RS-GLinCB is very sensitive to the parameter of $S$. This underperformance can be attributed to the fact that the warm-up period of RS-GLinCB is heavily dependent on the constant $S$ and $\kappa$ [Sawarni et al., 2024, Lemma 4.1].
@@
where $\beta_{t}(\delta)^{2}=\log\frac{1}{\delta}+\inf_{c_{t}\in(0,1]}\left\{d\log\frac{1}{c_{t}}+2SL_{t}c_{t}\right\}\leq\log\frac{1}{\delta}+d\log\left(e\vee\frac{2eSL_{t}}{d}\right)$. For Poisson bandits specifically, $L_{t}=e^{S}t+\log(d/\delta)$. This results in increasing cost in the optimization steps during the arm selection $X_{t}=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{X}_{t}}\max_{\theta\in\mathcal{C}_{t}(\delta)}\mathbf{x}^{\top}\theta$, as the algorithm needs to navigate a rapidly expanding nonconvex confidence region.
@@
E.3 Experiment on Forest Cover Type Data
@@
For this task, we set the horizon to $T=1000$ and the confidence parameter to $\delta=0.01$. After analyzing the data, we set $S=6$ and $\kappa=200$. We evaluated our algorithm against the same baselines used in the logistic bandit simulation experiment, running each method over 10 independent trials and averaging the results to report the regret and the running time. The error bars in the figures denote 99% confidence intervals for both regret and runtime.
Compared to synthetic data, real-world datasets exhibit higher noise and complexity, demanding careful exploration-exploitation trade-offs. Thus, we shrank the estimated confidence set of all the algorithms in a comparable way to achieve a better balance between exploration and exploitation in this real-world dataset. Traditional GLB algorithms are particularly sensitive to noise, often leading to excessive exploration and higher regret.